# An engine not a camera: Measuring performative power of online search

### v1

@@ Abstract
The power of digital platforms is at the center of major ongoing policy and regulatory efforts. To advance existing debates, we designed and executed an experiment to measure the power of online search providers, building on the recent definition of performative power. Instantiated in our setting, performative power quantifies the ability of a search engine to steer web traffic by rearranging results. To operationalize this definition we developed a browser extension that performs unassuming randomized experiments in the background. These randomized experiments emulate updates to the search algorithm and identify the causal effect of different content arrangements on clicks. We formally relate these causal effects to performative power. Analyzing tens of thousands of clicks, we discuss what our robust quantitative findings say about the power of online search engines. More broadly, we envision our work to serve as a blueprint for how performative power and online experiments can be integrated with future investigations into the economic power of digital platforms.
@@ 1 Introduction
At the heart of one of Europe’s most prominent antitrust case is a seemingly mundane question: How much can a search engine redirect traffic through content positioning? In 2017, the European Commission alleged that Google favored its own comparison shopping service by steering clicks away from search results towards Google’s own product comparison service. The technical centerpiece of the case was an ad-hoc data analysis about the position and display biases of Google search results. Google appealed the European Commission’s charges, pointing to, among other arguments, methodological errors. [^2] Case T-612/17, Google and Alphabet v Commission (Google Shopping), ECLI:EU:T:2021:763. The case is still pending, see Case C-48/22 P, Google and Alphabet v Commission (Google Shopping).
@@
The case is emblematic of a broader problem. Although urgently needed, there is currently no accepted technical framework for answering basic questions about the economic power of digital platforms. Lawyers, economists, and policy makers agree that traditional antitrust tools struggle with multi-sided platforms (Stigler Committee, 2019; Crémer et al., 2019). Among other issues, these tools are not able to capture behavioral aspects defining user-platform interactions. Against this backdrop, a recently developed concept, called performative power (Hardt et al., 2022), suggests a way to mitigate limitations of existing antitrust enforcement tools. Performative power measures how much a platform can causally influence platform users through its algorithmic actions. By directly relating power to a causal effect, it sidesteps the complexities underlying conventional market definitions and offers a promising framework to integrate data and experimental methods with digital market investigations. Although the definition of performative power enjoys appealing theoretical properties, a proof of its practical applicability was still missing.
@@
Figure 1: The ability to influence web traffic through content arrangement. Blue bars show average click probability observed for generic search results in position 1 to 6 on Google search under different counterfactual arrangements; default arrangement (left), swapping results 1 and 2 (middle), swapping results 1 and 3 (right). We see a significant effect of these arrangement changes on clicks. We will inspect this more formally in Section 5 where we also explore arrangement changes beyond reranking.
1.1 Our work
@@
In this work, we present a first proof of concept showing how to estimate performative power of a search engine in practice. The instantiation of performative power we are interested in concerns the ability of a search engine to impact web traffic through decisions of how to arrange content.
@@
Our core contribution is to design and implement an online experiment to provide quantitative insights into the causal effect of content arrangement on clicks for the two most widely used search engines, Google Search and Bing. Our experiment is based on a browser extension, called Powermeter, that emulates small updates to the platform’s algorithm by modifying how search results are displayed to users. We implement different counterfactual arrangements to inspect the effect of re-ranking and favored positioning (e.g., Ads or Shopping boxes) on clicks, both in isolation and jointly; the latter is relevant for the Google Shopping investigation. The arrangement to which a user is exposed is chosen at random every time they perform a search. We discuss several technical steps we implemented to take care of the internal validity of our experimental design.
@@
Using Powermeter we collected data of about 57,000 search queries from more than 80 different subjects, over the period of 5 months. The queries for any given user follow the distribution of queries under their every-day use of online search. Our data allows us to measure the causal effects of arrangement under natural interactions of users with the platform. Figure 1 provides a first glimpse into the observed effects. In summary, we find that consistently down-ranking the first element by one position causes a reduction in clicks of $42\%$ for the respective element on Google search on average. Down-ranking the same element by two positions yields a reduction of more than $50\%$. For Bing we find an even larger effect of ranking, although with less tight confidence intervals due to the small number of Bing queries performed by our participants. Inspecting different subsets of queries we find that the effect of position is larger for queries with a high number of candidate search results. When combining down-ranking with the addition of Ads or Shopping boxes, the effect of arrangement is even more pronounced, showing a distortion in click for the first result of $66\%$ averaged across queries where such elements are naturally present on Google search. To the best of our knowledge, we are the first to publish experimental insights into the display effects on Google search and Bing.
@@
Finally, we outline how to relate our measurements to performative power relevant in the context of digital market investigations. We take a recent antitrust investigation against Alphabet Inc. ran by the European Commission as a case study. Taken together, we hope our empirical and theoretical results can serve as a first blueprint for what future antitrust investigations of digital platforms’ market power might look like.
2 Background and related work
@@
Inspired by recent developments in the computer science literature, Hardt et al. (2022) proposed performative power as a measure of influence that predictive systems can have over participants. The term performativity more broadly refers to a causal effect of predictions on the target of prediction. The study of performativity in economics dates back to Morgenstern (1928); Grunberg and Modigliani (1954); Simon (1954). Performative power builds on the more recent development of performative prediction (Perdomo et al., 2020). Performative prediction introduces the concept of performativity to machine learning and exposes its role in data-driven predictions. We refer the reader to Hardt and Mendler-Dünner (2023) for a recent survey on the topic. A robust insight from performative prediction is that there are two ways to succeed at prediction. One is to learn patterns in the data-generating distribution. The other is to steer the data-generating distribution to a desirable state. Performative power (Hardt et al., 2022) builds on these insights and starts from a simple observation. The degree to which predictions are performative depends on the power—in terms of reach and scale—of the platform making the predictions. Thus, by measuring the degree to which predictions are performative, we can estimate the power of a platform. Formally, performative power relates the ability of a platform to steer the population of participants through predictions, to the causal effect of algorithmic actions. We provide the definition below.
@@
In words, performative power measures the ability of a firm to cause a change in the outcome on average across a population, by means of changing algorithmic actions. As such, performative power offers a family of definitions that can be instantiated flexibly in a given context. The distance notion could also be generalized, which we refrain from for the sake of presentation. For this work we discuss a concrete instantiation of performative power motivated by a recent antitrust investigation against Alphabet Inc. ran by the European Commission, [^3] European Commission, AT.39740, Google Search (Shopping), 27.06.2017. referred to as the Google Shopping case.
Google shopping case.
The subject of the investigation is the alleged conduct on Google search of introducing a) visually more appealing boxes for product comparisons on shopping queries in combination with b) an update to the ranking algorithm that leads to competing product comparison services appearing further down in the generic search results. The combination of these two conducts is made responsible for distorting outgoing web traffic from general search, and harm competition in the related specialised search markets (e.g. shopping). We are interested in providing empirical evidence into this effect.
We build on a result of Hardt et al. (2022) that formally relate performative power to the causal effect of position. For our investigation, we generalize the effect of position to encompassing algorithmic updates beyond reranking. These updates can be seen as changes to what is commonly referred to as choice architecture (Thaler and Sunstein, 2008) controlled by the platform’s algorithm. We focus on online search. This is worth emphasizing, as the causal effect of algorithmic changes is a result of users’ on-platform behavior and influenced by behavioral weaknesses, biases, and preferences. Thus, its strength strongly depends on the context in which users interact with the platform. To account for this we perform our experiments and measurements under natural interactions of users with the platform.
Causal effect of position.
Evidence for position bias was first reported in eye-tracking studies by Joachims et al. (2005); Guan and Cutrell (2007). Since then the importance of position bias in recommender systems more broadly, has been studied in multiple contexts. It play an important role when interpreting user feedback in recommender systems (Joachims et al., 2005), designing auctions (Aggarwal et al., 2006; Athey and Ellison, 2011), learning-to-rank (Agarwal et al., 2019; Qin et al., 2020; Chen et al., 2022; Zhang et al., 2023), debiasing click data (Craswell et al., 2008; Chuklin and and, 2015; Xv et al., 2022; Chen et al., 2023), and understanding incentives on different sides of the market (Ben-Porat and Tennenholtz, 2018; Hron et al., 2023; Hardt et al., 2023), to name a few. Most directly related to our work are efforts that aim to measure position bias: Ursu (2018) rely on public data collected under randomized result ordering to investigate the role of positioning on Expedia. Narayanan and Kalyanam (2015) investigated position bias in search advertising using a regression discontinuity design. Also focusing on online advertising, Agarwal et al. (2011) investigate position bias by experimentally randomizing bids to influence the ranking. With our work we demonstrate how a browser extension can serve as a useful tool for directly measuring the causal effect of position without having access to the platform’s algorithm. This allows us to provide experimental insights into multiple algorithmic changes. Browser extensions have previously been used for automatically collecting data to audit systems. Robertson et al. (2018) audit Google search for polarization on politically-related searches. Gleason et al. (2023) collect data via an extension to directly investigate the effect of search result components on clicks. Also the ongoing National Internet Observatory (NSF, 2021) relies on a browser extension to collect web traffic data. While most prior works focus on observational data, we use the extension to conduct online experiments.
@@ 3 Performativity in online search
We set up the causal question under investigation. Therefore, we model an online search platform as a distribution over events. An event is a triplet of a user query $Q$, content arrangement $A$ and click outcome $C$. A user query corresponds to a person visiting the search page and entering a search query in the search bar. The query is processed by the platform and results in an arrangement of content on the website. The mapping is typically defined by a proprietary pipeline involving a ranking algorithm that determines the order in which search results are ranked and displayed, including the positioning of components such as Ads or featured elements. Then, mediated by the arrangement, the user query leads to a click outcome $C$. The categorical random variable $C$ indexes the element clicked over by the user. It is a function of the user query and the arrangement.
@@ 3.1 The causal effect of arrangement
If clicks were solely determined by stable preferences, then we would see no effect. Performativity is the reason why we see an effect. Behavioral biases, weaknesses, the limited ability to process large amounts of data, and trust in the platform can be a source for performativity. The more performative the arrangement is, the stronger the effect. We use $a_{0}$ to refer to the reference arrangement of results on Google search. For a given user query $q$ we define the potential outcome of a click event under the arrangement $a$ as $C_{q}(a)$. The variable $C$ takes on categorical values, indexing the elements on the page. We mainly focus on clicks to general search results. Let $\{c_{1},...c_{K}\}$ denote the top $K$ general search results indexed in the order they appear under $a_{0}$.
@@
$\mathrm{\delta}^{i}(a^{\prime})=\operatorname*{\mathbb{E}}\big{[}1\{C_{q}(a^{\prime})=c_{i}\}\big{]}-\operatorname*{\mathbb{E}}\big{[}1\{C_{q}(a_{0})=c_{i}\}\big{]}\,,$ (2)
where expectations are taken over $q\in\mathcal{Q}$ and the randomness in the potential outcome.
In words, the performativity gap quantifies how much the expected click through rate of search item $c_{i}$ on average across queries $\mathcal{Q}$ changed, had the platform deployed arrangement $a^{\prime}$ instead of arrangement $a_{0}$. We are particularly interested in counterfactuals $a^{\prime}$ that constitute small changes to the original arrangement $a_{0}$, such as the slight reordering of search results, or the addition of visually appealing elements on the website. The performativity gap offers a generalization of the causal effect of position defined in Hardt et al. (2022), to cover algorithmic updates beyond reranking. We will discuss concrete instantiations in Section 4.2.
More broadly, the effect of small arrangement changes provides a lower bound on the ability of a platform to impact clicks with potentially more significant algorithmic changes. The following result generalizing Theorem 8 in Hardt et al. (2022) is a direct consequence:
Theorem 1 (Generalization of Hardt et al. (2022); Theorem 8).
@@
This quantity serves as a way to denote the fraction of clicks taken away from content item $c_{i}$ as a result of the update $a_{0}\rightarrow a^{\prime}$. Also, it helps to express power relative to a base click through rate as it offers a more interpretable quantity for investigators.
3.2 Estimation using an RCT
To estimate the performativity gap, we rely on a randomized controlled trial (RCT), the gold standard methodology to estimate causal effects (Fisher, 1935; Imbens and Rubin, 2015). Given the fundamental challenge of causal inference that we can not observe a search query simultaneously exposed to different arrangements, the idea of an RCT is to randomly select, for each query $q\in\mathcal{Q}$, the arrangement they are exposed to. We write $Q_{a}\subset\mathcal{Q}$ for the subset of queries that are exposed to treatment $A=a$. We also refer these subsets as treatment groups. Observing the click events for each queries allows us to obtain an estimate of the performativity gap as
@@ 3.2 Estimation using an RCT
For this ${\mathrm{\bar{\delta}}}^{i}(a^{\prime})$ to provide an unbiased estimate of $\delta^{i}(a^{\prime})$, we rely on the following application of the stable unit treatment value assumption (SUTVA) (Rubin, 1980) to be satisfied:
@@
In the following section we discuss steps we take in our experimental design towards justifying this Assumption 1. Beyond requiring relatively weak assumptions, the advantage of using an experimental approach to measure the performativity gap is that, while the mechanism mapping user queries to clicks can be arbitrarily complex, this complexity does not affect the experiment. Aspects such as users’ preference for clicking links on the left side of the screen (Dimitrov et al., 2016), users’ trust in the platform ranking (Pan et al., 2007), the effect of visually appealing elements (Fubel et al., 2023), and other behavioral biases will naturally enter our measurement. We directly measure their effect on the relevant quantity of interest, namely, the resulting clicks.

### v2

@@ Abstract
The power of digital platforms is at the center of major ongoing policy and regulatory efforts. To advance existing debates, we designed and executed an experiment to measure the performative power of online search providers. Instantiated in our setting, performative power quantifies the ability of a search engine to steer web traffic by rearranging results. To operationalize this definition we developed a browser extension that performs unassuming randomized experiments in the background. These randomized experiments emulate updates to the search algorithm and identify the causal effect of different content arrangements on clicks. Analyzing tens of thousands of clicks, we discuss what our robust quantitative findings say about the power of online search engines, using the Google Shopping antitrust investigation as a case study. More broadly, we envision our work to serve as a blueprint for how the recent definition of performative power can help integrate quantitative insights from online experiments with future investigations into the economic power of digital platforms.
@@ 1 Introduction
At the heart of one of Europe’s most prominent antitrust case is a seemingly mundane question: How much can a search engine redirect traffic through content positioning? In 2017, the European Commission alleged that Google favored its own comparison shopping service by steering clicks away from search results towards Google’s own product comparison service. The technical centerpiece of the case was an ad-hoc data analysis about the position and display biases of Google search results. Google appealed the European Commission’s charges, pointing to, among other arguments, methodological errors. [^2] Case C-48/22 P, Google and Alphabet v Commission (Google Shopping), ECLI:EU:C:2024:14.
@@
The case is emblematic of a broader problem. Although urgently needed, there is currently no accepted technical framework for answering basic questions about the economic power of digital platforms. Lawyers, economists, and policy makers agree that traditional antitrust tools struggle with multi-sided platforms (Stigler Committee, 2019; Crémer et al., 2019). Against this backdrop, a recently developed concept from the machine learning literature, called performative power (Hardt et al., 2022), suggests a way to mitigate limitations of existing antitrust enforcement tools. Performative power measures how much a platform can causally influence platform users through its algorithmic actions. By directly relating power to a causal effect, it sidesteps the complexities underlying conventional market definitions and offers a promising framework to integrate data and experimental methods with digital market investigations. Although the definition of performative power enjoys appealing theoretical properties, a proof of its practical applicability was still missing.
@@
Figure 1: The ability to influence web traffic through content arrangement. Blue bars show average click probability observed for generic search results in position 1 to 6 on Google search under different counterfactual arrangements; default arrangement (left), swapping results 1 and 2 (middle), swapping results 1 and 3 (right). We provide a detailed discussion in Section 5 where we also explore arrangement changes beyond reranking.
1.1 Our contributions
@@
In this work we present a first proof of concept showing how to use performative power as an investigative tool in practice. The instantiation of performative power we consider is motivated by the recent Google Shopping antitrust investigation ran by the European Commission against Alphabet Inc. It concerns the ability of a search engine to impact web traffic through decisions of how to arrange content.
@@
Our core contribution is to design and implement an online experiment to establish a lower bound on performative power by providing quantitative insights into the causal effect of algorithmic updates on clicks for the two most widely used search engines, Google Search and Bing. Our experiment is based on a browser extension, called Powermeter, that emulates updates to the platform’s algorithm by modifying how search results are displayed to users. We implement different counterfactual arrangements to inspect the effect of re-ranking and favored positioning (e.g., Ads or Shopping boxes) on clicks, both in isolation and jointly. The arrangement to which a user is exposed is chosen at random every time they perform a search. We discuss several technical steps we implemented to take care of the internal validity of our experimental design.
@@
Using Powermeter we collected data of about 57,000 search queries from more than 80 different subjects, over the period of 5 months. The queries for any given user follow the distribution of queries under their every-day use of online search. Figure 1 provides a first glimpse into the observed effects of arrangement on clicks. In summary, we find that consistently down-ranking the first element by one position causes an average reduction in clicks of $42\%$ for the respective element on Google search. Down-ranking the same element by two positions yields a reduction of more than $50\%$. The causal effect is illustrated by the change in the hight of the bars. For Bing we find an even larger effect of ranking, although with less tight confidence intervals due to the small number of Bing queries performed by our participants. When combining down-ranking with the addition of Ads or Shopping boxes, the effect of arrangement is even more pronounced, showing a distortion in clicks for the first result of $66\%$ averaged across queries where such elements are naturally present on Google search. Inspecting different subsets of queries we find that the effect of position is larger for queries with a high number of candidate search results. To the best of our knowledge, we are the first to offer independent quantitative experimental insights into display effects on Google search and Bing.
@@
Finally, we outline how to formally relate our quantitative piece of evidence to questions about self-preferencing relevant in the context of the Google Shopping case. Taken together, we hope our empirical and theoretical results can serve as a first blueprint for what future antitrust investigations of digital platforms’ market power based on performative power might look like.
2 Preliminaries and related work
@@
The concept of performative power offers an alternative notion of power inspired by recent developments in performative prediction (Perdomo et al., 2020) within the computer science literature. We refer the reader to Hardt and Mendler-Dünner (2023) for a recent survey on the topic. A robust insight from performative prediction is that beyond learning patterns in data, the ability to steer the data-generating distribution similarly factors into a predictive system’s performance. Performative power (Hardt et al., 2022) recognizes that the ability to steer depends on power—in terms of reach and scale—of the platform making the predictions. Thus, the core idea behind performative power is to measure the degree to which predictions are performative to obtain an estimate of the power of a platform. Formally, performative power relates the ability of a platform to steer a population of participants, to the causal effect of algorithmic actions.
@@
Performative power is a measure of influence that predictive systems can have over their participants. It offers a family of definitions that can be instantiated flexibly in a given context. The specific meaning is determined by each instantiation. Performative power can be applied forward-looking to understand whether a platform has the ability to plausibly cause a specific change, as well as in retrospect to measure the effect of an observed conduct. In this work we use performative power to quantify the effect of an algorithmic update $a^{*}$ central to a recent antitrust investigation against Alphabet Inc. ran by the European Commission. [^3] European Commission, AT.39740, Google Search (Shopping), 27.06.2017. For this purpose, we instantiate $\mathcal{A}$ with a set of conservative and implementable counterfactuals such as to provide a plausible lower bound on the effect of $a^{*}$.
The Google Shopping case.
[…]
Display effects and behavioral biases.
Consumer choices on digital platforms are critically mediated by how platforms present content to users. Choice architecture designs (Thaler and Sunstein, 2008), presentation bias (Yue et al., 2010), position bias (Joachims et al., 2005; Guan and Cutrell, 2007), and trust bias (Pan et al., 2007) are known to play an important role. There is a rich literature in machine learning aiming to mechanistically understand such biases for debiasing click data (Craswell et al., 2008; Chuklin and and, 2015; Xv et al., 2022; Chen et al., 2023; Agarwal et al., 2019; Qin et al., 2020; Chen et al., 2022; Zhang et al., 2023), building better ranking models and auctions (Aggarwal et al., 2006; Athey and Ellison, 2011), and interpreting user feedback in recommender systems (Joachims et al., 2005), to name a few. Unfortunately behavioral aspects often resist a clean mathematical specification. By focusing on measuring a directly observable statistic, performative power circumvents the challenges of modeling behavioral biases for monitoring, auditing and measuring digital economies.
Measuring the effect of algorithmic updates.
Several prior works have been interested in measuring the effects of potential arrangement changes on consumption on online platforms. For example, Ursu (2018) rely on public data collected under randomized result ordering to investigate the role of positioning on Expedia. Narayanan and Kalyanam (2015) investigated position bias in search advertising using a regression discontinuity design. Also focusing on online advertising, Agarwal et al. (2011) investigate position bias by experimentally randomizing bids to indirectly influence the ranking. Similarly in information retrieval researchers have studied active interventions in the form of order randomization (Radlinski and Joachims, 2006), or relied on harvesting click data collected under multiple historical rankings (Fang et al., 2019). In our work we collect experimental data ourselves. We use a browser extension to emulate the algorithmic updates of interest without requiring control over the platform’s algorithm.
Browser extensions have previously been used as a tool for automatically collecting data to audit systems. Robertson et al. (2018) audit Google search for polarization on politically-related searches. Gleason et al. (2023) collect data via an extension to directly investigate the effect of search result components on clicks. Also the ongoing National Internet Observatory (NSF, 2021) relies on a browser extension to collect web traffic data. While prior works focus on collecting observational data for monitoring systems, we use the extension to conduct online experiments.
@@ 3 Performativity in online search
We start by formalizing the causal question under investigation. We model an online search platform as a distribution over events. An event is a triplet of a user query $Q$, content arrangement $A$ and click outcome $C$. A user query corresponds to a person visiting the search page and entering a search query in the search bar. The query is processed by the platform and results in an arrangement of content on the website. The mapping is typically defined by a proprietary pipeline involving a ranking algorithm that determines the order in which search results are ranked and displayed, including the positioning of components such as Ads or featured elements. Then, mediated by the arrangement, the user query leads to a click outcome $C$. The categorical random variable $C$ indexes the element clicked over by the user. It is a function of the user query and the arrangement.
@@ 3.1 The causal effect of arrangement
If clicks were solely determined by stable preferences, then we would see no effect. Performativity is the reason why we see an effect. Display baises, the limited ability to process large amounts of data, and trust in the platform can be a source for performativity. The more performative the arrangement is, the stronger the effect. We use $a_{0}$ to refer to the reference arrangement of results on Google search. For a given user query $q$ we define the potential outcome of a click event under the arrangement $a$ as $C_{q}(a)$. The variable $C$ takes on categorical values, indexing the elements on the page. Let $\{c_{1},...c_{K}\}$ denote the top $K$ general search results indexed in the order they appear under $a_{0}$.
@@
$\mathrm{\delta}^{i}(a^{\prime})=\operatorname*{\mathbb{E}}\big{[}1\{C_{q}(a^{\prime})=c_{i}\}\big{]}-\operatorname*{\mathbb{E}}\big{[}1\{C_{q}(a_{0})=c_{i}\}\big{]}\,,$
where expectations are taken over queries $q\in\mathcal{Q}$ and the randomness in the potential outcome.
The performativity gap quantifies how much the click through rate of search item $c_{i}$ changed, in expectation across queries $\mathcal{Q}$, had the platform deployed arrangement $a^{\prime}$ instead of arrangement $a_{0}$. The following result generalizes Theorem 8 in Hardt et al. (2022):
Theorem 1 (Lowerbound on performative power).
@@
Note that the instantiation of performative power in Theorem 1 to which we relate the performativity gap measures a platform’s ability to steer outgoing traffic from its online search website. We will discuss how to relate this notion to a broader discussion of the power of online search in vertically integrated markets in Section 6.
@@
This quantity serves as a way to denote the fraction of clicks taken away from content item $c_{i}$ as a result of the update $a_{0}\rightarrow a^{\prime}$. Also, as we will see, it helps to express performative power relative to a base click through rate which offers a more interpretable quantity for investigators.
3.2 Estimating the performativity gap using an RCT
To estimate the performativity gap for the individual arrangements, we rely on a randomized controlled trial (RCT), the gold standard methodology to estimate causal effects (Fisher, 1935; Kohavi et al., 2009; Imbens and Rubin, 2015). As we can not observe a search query simultaneously exposed to different arrangements, the idea of an RCT is to randomly select, for each query $q\in\mathcal{Q}$, the arrangement they are exposed to. We write $Q_{a}\subset\mathcal{Q}$ for the subset of queries that are exposed to treatment $A=a$. We also refer to these subsets as treatment groups. Comparing the click events across groups allows us to obtain an estimate of the performativity gap as
@@ 3.2 Estimating the performativity gap using an RCT
For ${\mathrm{\bar{\delta}}}^{i}(a^{\prime})$ to provide an unbiased estimate of $\delta^{i}(a^{\prime})$, we rely on an application of the stable unit treatment value assumption (SUTVA) (Rubin, 1980), referred to as isolation assumption by Bottou et al. (2013):
@@
This assumption justifies why we can interleave the measurement of different interventions in our experiment. It implies that the intervention performed on one query does not change individuals’ browsing behavior in response to subsequent queries. For this to reasonably hold, individual interventions under should not impede user experience in a lasting manner. In the following we discuss steps we take in our experimental design towards justifying Assumption 1.
More broadly, the key advantage of using an experimental approach to measure the performativity gap is that, while the mechanism mapping user queries to clicks can be arbitrarily complex, this complexity does not affect the experiment. Aspects such as users’ preference for clicking links on the left side of the screen (Dimitrov et al., 2016), the effect of visually appealing elements (Fubel et al., 2023), users’ trust in the platform (Pan et al., 2007), or the relevance gap between search results will naturally enter our measurement.